trying to help my 7th grader , with his 8th grerade pre algebra. He knows the basis of this problem, but isn't sure what to do since the x is in a fraction.
6(x/2-4) = 21+ 6x

Answer by Alan3354(69443)

6(x/2-4) = 21+ 6x
6*(x/2) = 3x
------
6(x/2-4) = 21+ 6x
3x-24 = 6x + 21
3x = 6x + 45
-3x = 45
x = -15